# Thermally-induced mimicry of quantum cluster excitations and implications for the magnetic transition in FePX3

### v1

@@ VI Discussion
One prominent family of systems to which the local cluster model has been applied is the Cr3+ spinels, in which $S=3/2$ ions reside on a frustrated pyrochlore lattice. Above a spin-Peierls transition, rings of scattering intensity are observed in momentum space, which have been interpreted as hexagonal loop excitations in real space. Constant energy slices at higher energies display qualitatively different momentum dependence ascribed to different modes of tetramer excitations [39, 41]. In MgCr2O4, excitations below the magnetic ordering temperature show a flat mode at low energy, which has been suggested as evidence of the persistence of cluster fluctuations down to low temperature. Whilst these models successfully describe the momentum dependence at a given energy, they do not describe the energy-resolved spectra. In Ref.  52, some of us, along with coauthors, demonstrated that above $T_{N}$, the spectra could be understood semiclassically in terms of harmonic spin wave excitations above the disordered ground state associated with the macroscopically-degenerate classical spin liquid phase. Further to this, in Ref. 53, the low energy resonance mode was shown to be dispersive, consistent with coherent spin wave excitations.
@@
Thus, the mechanism underlying the broadened response in FePX3 differs from the case of a highly frustrated magnet such as MgCr2O4. In FePX3, the broadening originates from classical nonlinearities of the spin dynamics induced by thermal fluctuations. In MgCr2O4 the broadened response measured just above $T_{N}$ is primarily a result of harmonic fluctuations averaged over a quasi-degenerate manifold of spin configurations. This difference reflects the distinct origins of the spin disorder in the two systems. In MgCr2O4, $T_{N}$ is small compared to the exchange interaction scale ($\Theta_{W}/T_{N}\approx 30$[54] and hence the nonlinear terms in the equation of motion are small at or just above $T_{N}$. In contrast, $T_{N}$ is roughly on the same order of the Weiss constant [55] for FePX3. It is, therefore, to be expected that the nonlinear fluctuations close to $T_{N}$ are sizeable and play a key role in the dynamics of the latter compound.
@@
The results presented in this paper suggest that the finite temperature dynamics in FePX3 can be well described using semiclassical methods. In this picture, the order-disorder phase transition at $T_{N}$ can be viewed as the overcoming of the anisotropy gap, which protects the long-range magnetic ordering in two dimensions. The sharp nature of the transition makes the experimental observation of the gap closure challenging since the gap rapidly closes as thermal fluctuations overcome the anisotropy energy. Similarities can be observed between the order-disorder transition in FePX3 and CrI3, a weakly anisotropic ferromagnet for which the phase transition is weakly first order and the gap goes to zero at the transition whilst the spin stiffness remains finite [56].
@@
The finite spin stiffness at the phase transition also accounts for the highly structured magnetic fluctuations observed above $T_{N}$, which arise from strong short-range correlations. The nature of the gap destruction in FePX3 is somewhat different from the traditional picture of gap closure such as the field induced transition of a quantum antiferromagnet in field, where the singlet condensate is replaced by a magnon condensate as the magnon gap closes. Here, the magnon density is the order parameter, growing continuously as the field increases. In FePX3, thermal fluctuations protruding from the antiferromagnetic wavevector overcome the anisotropy energy abruptly as the temperature increases above $T_{N}$. In this sense, we can view this as a closure of the anisotropy gap from below.
@@
We note that the first-order nature of the transition, along with the observation of a structural distortion in FePS3 [57] are suggestive of a magnetoelastic origin for the phase transition in FePS3. To our knowledge, no evidence of such distortion has been observed in FePSe3 [58], and the success of our model in describing the fluctuations at, and above, $T_{N}$ suggests that magnetoelasticity may not drive the order-disorder transition. Measuring the magnetic fluctuations close to the ordering transition in FePS3 would be interesting to compare with its sister compound.
@@ VII Conclusion
In this paper, we have presented semiclassical calculations describing the magnetic order and dynamics of the 2D van der Waals antiferromagnetic systems FePX3. Using a combination of linear spin wave theory and Landau-Lifshitz dynamics for dipolar spins, we have reproduced the zero-temperature spin wave excitations and the short-range fluctuations at finite temperature with a common model. Our simulations point to the overcoming of the anisotropy energy by thermally-induced antiferromagnetic fluctuations as the origin of the first-order magnetic order-disorder phase transition. We speculate that these results are general to a broad range of large-$S$ 2D van der Waals magnets for which a strong easy anisotropy sets the magnetic ordering temperature. These results further motivate careful temperature-dependent spectroscopic and diffraction measurements in FePX3.
@@ Acknowledgements.
H.L was funded by a Research Fellowship from the Royal Commission for the Exhibition of 1851. M.M acknowledges support from U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division under award DE-SC-0018660. H.L acknowledges useful discussions with A. R. Wildes, C. D. Batista, D. Dahlbom, K. Barros and X. Bai.
Appendix A Renormalized Classical Spin (RCS) theory
In order to compare the calculations in the literature with our work, we rescale the Hamiltonian parameters to account for the corrections to leading order in $1/S$ according to Ref. 45. The correction to the single ion anisotropy is
@@ Appendix A Renormalized Classical Spin (RCS) theory
Appendix B Decorrelation
@@
Appendix C Determining $T_{N}^{\mathrm{cl}}$
We calculate the classical ordering temperature, $T_{N}^{\mathrm{cl}}$ for each model under consideration by integrating about the Bragg peak as a function of temperature and numerically differentiating the order parameter as a function of temperature. The site of the discontinuity was taken to be the classical ordering temperature $T_{N}^{\mathrm{cl}}$. Calculations were performed on an (24,24,2) supercell with 1000 samples at each of 59 temperatures, spaced exponentially from both directions to provide dense coverage around the transition temperature.

### v3

@@ VI Discussion
One prominent family of systems to which the local cluster model has been applied is the Cr3+ spinels, in which $S=3/2$ ions reside on a frustrated pyrochlore lattice. Above a spin-Peierls transition, rings of scattering intensity are observed in momentum space, which have been interpreted as hexagonal loop excitations in real space. Constant energy slices at higher energies display qualitatively different momentum dependence ascribed to different modes of tetramer excitations [39, 41]. In MgCr2O4, excitations below the magnetic ordering temperature show a flat mode at low energy, which has been suggested as evidence of the persistence of cluster fluctuations down to low temperature. Whilst these models successfully describe the momentum dependence at a given energy, they do not describe the energy-resolved spectra. In Ref.  68, some of us, along with coauthors, demonstrated that above $T_{N}$, the spectra could be understood semiclassically in terms of harmonic spin wave excitations above the disordered ground state associated with the macroscopically-degenerate classical spin liquid phase. Further to this, in Ref. 69, the low energy resonance mode was shown to be dispersive, consistent with coherent spin wave excitations.
@@
Thus, the mechanism underlying the broadened response in FePSe3 differs from the case of a highly frustrated magnet such as MgCr2O4. In FePSe3, the broadening originates from classical nonlinearities of the spin dynamics induced by thermal fluctuations. In MgCr2O4 the broadened response measured just above $T_{N}$ is primarily a result of harmonic fluctuations averaged over a quasi-degenerate manifold of spin configurations. This difference reflects the distinct origins of the spin disorder in the two systems. In MgCr2O4, $T_{N}$ is small compared to the exchange interaction scale ($\Theta_{W}/T_{N}\approx 30$[70] and hence the nonlinear terms in the equation of motion are small at or just above $T_{N}$. In contrast, $T_{N}$ is roughly on the same order of the Weiss constant [71] for FePX3. It is, therefore, to be expected that the nonlinear fluctuations close to $T_{N}$ are sizeable and play a key role in the dynamics of the latter compound.
@@
The results presented in this paper suggest that the finite temperature dynamics in FePSe3 can be well described using semiclassical methods. In this picture, the order-disorder phase transition at $T_{N}$ can be viewed as the overcoming of the anisotropy gap, which protects the long-range magnetic ordering in two dimensions. The sharp nature of the transition makes the experimental observation of the gap closure challenging since the gap rapidly closes as thermal fluctuations overcome the anisotropy energy. Similarities can be observed between the order-disorder transition in FePSe3 and CrI3, a weakly anisotropic ferromagnet for which the phase transition is weakly first order and the gap goes to zero at the transition whilst the spin stiffness remains finite [65].
@@
The finite spin stiffness at the phase transition also accounts for the highly structured magnetic fluctuations observed above $T_{N}$, which arise from strong short-range correlations. The nature of the gap destruction in FePSe3 is somewhat different from the traditional picture of gap closure such as the field induced transition of a quantum antiferromagnet in field, where the singlet condensate is replaced by a magnon condensate as the magnon gap closes. Here, the magnon density is the order parameter, growing continuously as the field increases. In FePSe3, thermal fluctuations protruding from the antiferromagnetic wavevector overcome the anisotropy energy abruptly as the temperature increases above $T_{N}$. In this sense, we can view this as a closure of the anisotropy gap from below.
@@
We note that the first-order nature of the transition, along with the observation of a structural distortion in FePS3 [45] are suggestive of a magnetoelastic origin for the phase transition in FePS3. To our knowledge, no evidence of such distortion has been observed in FePSe3 [72], and the success of our model in describing the fluctuations at, and above, $T_{N}$ suggests that magnetoelasticity may not drive the order-disorder transition. Measuring the magnetic fluctuations close to the ordering transition in FePS3 would be interesting to compare with FePSe3. The situation in FePSe3 appears reminiscent of FeI2 which exhibits a first order thermodynamic phase transition without the observation of a structural transition [73]. There are further similarities between the dynamics in FePSe3 and FeI2, where the spectrum remains gapped as $T_{N}$ is approached [52].
@@ VII Conclusion
In this paper, we have presented semiclassical calculations describing the magnetic order and dynamics of the 2D van der Waals antiferromagnetic systems FePSe3. Using a combination of linear spin wave theory and Landau-Lifshitz dynamics for dipolar spins, we have reproduced the zero-temperature spin wave excitations and the short-range fluctuations at finite temperature with a common model. Our simulations point to the overcoming of the anisotropy energy by thermally-induced antiferromagnetic fluctuations as the origin of the first-order magnetic order-disorder phase transition. We speculate that these results are general to a broad range of large-$S$ 2D van der Waals magnets for which a strong easy anisotropy sets the magnetic ordering temperature. These results further motivate careful temperature-dependent spectroscopic and diffraction measurements in FePSe3.
@@ Acknowledgements.
H.L was funded by a Research Fellowship from the Royal Commission for the Exhibition of 1851. M.M acknowledges support from U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division under award DE-SC-0018660. H.L acknowledges useful discussions with L. Chen, J.-H. Chung, R. J. Birgeneau, P. Dai, A. R. Wildes, C. D. Batista, D. Dahlbom, K. Barros and X. Bai. The research data supporting this publication can be accessed at http://dx.doi.org/10.6084/m9.figshare.29077010 [74].
Appendix A FePS3
Table 2: Exchange parameters (in meV) adapted from Wildes et al. [31] using the RCS to be consistent with conventions used in calculations for FePSe3 in this paper. A small interchain coupling has been added to reproduce the reported magnetic ground state. Exchange couplings from include a factor of two to account for the difference in Hamiltonian definition.
$J_{1}$2.02
$J_{2}$  0.30
$J_{3}$  0.9
$J_{c}$  0.01
$K_{1}$0.20
$K_{2}$0.04
$D$  3.36
The crystal structures of FePS3 differs from FePSe3 having a broken $C_{3}$ symmetry and crystallizing in the monoclinic $C2/m$ space group. The broken $C_{3}$ symmetry splits the nearest-neighbor bonds on the honeycomb lattice into inequivalent zig-zag, $J_{1a}$ and leg, $J_{1b}$ bonds. The subtle nature of the distortion has nonetheless motivated models where both $J_{1a}=J_{1b}$ and $J_{1a}\neq J_{1b}$ in FePS3 [32, 31], however, the most successful fits have employed models where $J_{1a}=J_{1b}$.
A $J_{1}$-$J_{2}$-$J_{3}$-$K_{1}$-$K_{2}$ model with single-ion anisotropy best describes the inelastic neutron scattering data [31], with parameters qualitatively similar to the fitted values for FePSe3 [1] including with a sizable ferromagnetic nearest-neighbor and antiferromagnetic next-nearest-neighbor Heisenberg exchange, a strong uniaxial anisotropy, and a ferromagnetic biquadratic nearest-neighbor exchange. For completeness, we quote the fitted parameters for FePS3 in Table 2.
Appendix B Renormalized Classical Spin Theory
In order to compare the calculations in the literature with our work, we rescale the Hamiltonian parameters to account for the corrections to leading order in $1/S$ according to Ref. 46. The correction to the single ion anisotropy is
@@ Appendix B Renormalized Classical Spin Theory
Appendix C Landau Lifshitz Dynamics
The finite temperature dynamics are simulated the Landau Lifshitz equations with noise and damping terms that are chosen to enforce the fluctuation-dissipation theorem. In this appendix, we reproduce a number of the key equations for completeness.
The Landau Lifshitz equation, for the three component spin dipole, $\mathbf{S}_{j}$ with damping and noise is given by
$\frac{d\mathbf{S}_{j}}{dt}=-\mathbf{S}_{j}\times\left(\boldsymbol{\xi}_{j}+\frac{\partial\mathcal{H}}{\partial\mathbf{S}_{j}}-\lambda\mathbf{S}_{j}\times\frac{\partial\mathcal{H}}{\partial\mathbf{S}_{j}}\right).$ (4)
The noise term is chosen to be a Gaussian white noise with
$\langle\xi_{j}^{\alpha}(t)\rangle=0$ (5)
$\langle\xi_{i}^{\alpha}(t)\xi_{j}^{\beta}(t^{\prime})\rangle=2\lambda k_{B}T\delta_{ij}\delta_{\alpha\beta}\delta(t-t^{\prime}).$ (6)
Equation 4 was numerically integrated using the explicit Heun method in the Sunny.jl spin dynamics package [48]. Interested readers are referred to Refs. [49, 50, 48] for further details.
Appendix D Decorrelation
@@
Appendix E Determining $T_{N}^{\mathrm{cl}}$
We calculate the classical ordering temperature, $T_{N}^{\mathrm{cl}}$ for each model under consideration by integrating about the Bragg peak as a function of temperature and numerically differentiating the order parameter as a function of temperature. The site of the discontinuity was taken to be the classical ordering temperature $T_{N}^{\mathrm{cl}}$. Calculations were performed on an $24\times 24\times 2$ supercell with 1000 samples at each of 59 temperatures, spaced exponentially from both directions to provide dense coverage around the transition temperature.